# Autonomous Evaluation and Refinement of Digital Agents

### v2

@@ Abstract
We show that domain-general automatic evaluators can significantly improve the performance of agents for web navigation and device control. We experiment with multiple evaluation models that trade off between inference cost, modularity of design, and accuracy. We validate the performance of these models in several popular benchmarks for digital agents, finding between 74.4 and 92.9% agreement with oracle evaluation metrics. Finally, we use these evaluators to improve the performance of existing agents via fine-tuning and inference-time guidance. Without any additional supervision, we improve state-of-the-art performance by 29% on the popular benchmark WebArena, and achieve a 75% relative improvement in a challenging domain transfer scenario. We release our code and data at https://github.com/Berkeley-NLP/Agent-Eval-Refine.
@@ 1 Introduction
We first evaluate our proposed approach on its ability to match oracle evaluation metrics using WebArena (Zhou et al., 2024) and Android-in-the-Wild (AitW; Rawles et al., 2023), achieving accuracies up to 82.1 and 92.9% respectively. We then show how these evaluation models can be used to refine existing agents through inference-time guidance or during training, without access to any hand-designed evaluation functions or additional demonstration data (Figure 1, right). When integrated as the reward function in Reflexion (Shinn et al., 2023), the evaluator enhances the best-performing GPT-4 WebArena agent’s success rate by up to 29% of relative improvement. Additionally, we evaluate in a domain transfer setting to device control in iOS, for which there is no existing benchmark environment or training data. When using our evaluation models to filter sampled trajectories to be used in behavior cloning, we see relative improvements of 75% accuracy in this domain.
@@ 2 Related Work
Early modeling of language-conditioned autonomous agents focused on approaches that include semantic parsing (Allen et al., 2007; Xu et al., 2021; Li et al., 2020), reinforcement learning (Branavan et al., 2009; 2010), and imitation learning (Humphreys et al., 2022). The strength of pretrained language and language-and-vision modeling has renewed interest in building language-conditioned digital agents (Zhang & Zhang, 2023; Hong et al., 2023; Zhou et al., 2024; Deng et al., 2023; Wang et al., 2023a; Gur et al., 2024). For example, baseline approaches to WebArena (Zhou et al., 2024) use few-shot prompting with language-only models, representing the environment state and action space with its document object model (DOM). More recent works in building these agents have moved from language-only modeling to vision-language modeling, representing the environment state space as its rendered pixel representation instead of relying on a DOM. Another line of work has applied inference-time techniques to improve model’s performance, for example with inference-time exploration (Zhang et al., 2023), intermediate plan revision (Zhang et al., 2024) and error correction (Wang et al., 2024), and self-critique (Wu et al., 2024) on GPT-4 or GPT-4V. Concurrent to our work, OS-Copilot (Wu et al., 2024) proposes a self-critique component to autonomously refine Mac device control agents, implementing the critic as a LM that reasons about proposed tool implementations and error messages. In contrast to our work, this critic does not evaluate actual agent behavior in the execution environment or used in model training.
@@
Recently, there has been renewed interest in methods for improving policies at training (Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Abdulhai et al., 2023) or inference (Shinn et al., 2023; Yao et al., 2023; Wu et al., 2024) time without human supervision. Approaches like Reflexion (Shinn et al., 2023) assume access to an external evaluation function, leveraging it as the supervision signal to guide policy improvement at inference time. In contrast, we study applications of inference-time autonomous refinement without requiring access to the external evaluation function, and show that using our proposed domain-general evaluation models improves agent success rate by 29%. Meanwhile, methods which bootstrap policies with supervised training and refine them with reinforcement learning have been widely adopted; our proposed evaluators enable this paradigm in open, realistic digital agent scenarios, providing a relative improvement in performance of over 70%. Concurrent to our work, WebVoyager (He et al., 2024) also explores using GPT-4V as an automated proxy for human evaluation of web agents, though is not the primary focus of their work, and neither performs in-depth analysis of the quality of its judgments, nor explores its applicability to improving agents.
@@
In parallel, the community has focused on demonstration-based benchmarks that do not include an executable simulation environment, including PIXELHELP (Li et al., 2020), MoTIF (Burns et al., 2022), Mind2Web (Deng et al., 2023), and AitW (Rawles et al., 2023). Notably, Mind2Web and AitW contain over 2K and 715K human trajectories respectively on a wide range of web navigation and device control tasks. Though primarily used for model training (Rawles et al., 2023; Deng et al., 2023; Hong et al., 2023; Zhang & Zhang, 2023), these datasets are also used for evaluating digital agents through reference-based metrics like action matching score. In this setting, an agent is given the prefix of a human demonstration and evaluated on its prediction of the next action to take. However, this metric requires human demonstrations and does not directly reflect agent’s performance in real-world because it does not account for consequences of an agent’s sequential decision process, alternative actions that diverge from the demonstration.
@@ 3.2 Modular Caption-then-Reason Approach
To improve a captioner’s ability to provide detailed, well-formatted descriptions, we collect a dataset of screenshots paired with descriptions, and use it to fine-tune an open-weight VLM. We first acquire screenshots from a variety web and device control domains, then use GPT-4V to provide an initial detailed description for each screenshot. We manually filter out or fix apparent errors in GPT-4V’s output, resulting a total of 1,263 data points. [^3] Table 3 in Section A.2 contains details of data sources and sizes. We use this data to fine-tune the QWen-VL (Bai et al., 2023) model. During both finetuning and at inference time, we provide text recognition results from EasyOCR [^4] https://github.com/JaidedAI/EasyOCR as an additional input to the VLM to reduce hallucination.
@@ 4 Experiments and Results
Our rationale behind experiment design is to cover a broad range of domains and challenges. We use WebArena for both evaluation and inference-time refinement, as its built-in evaluation functions facilitate direct comparison. Android-in-the-Wild (AitW) is chosen for evaluation since it is widely used for training and evaluating Android agents, and is typically evaluated using a reference-based metric instead of task success. Lastly, we refine a model through filtered behavior cloning on iOS, where data scarcity poses a significant challenge to supervised methods.
@@
  • Captioner + Mixtral: Modular approach (Section 3.2) using a finetuned QWen-VL (Bai et al., 2023) to generate a trajectory description, and Mixtral (Jiang et al., 2024) to provide the final evaluation.
  • Captioner + GPT-4: Modular approach (Section 3.2) using a finetuned QWen-VL to generate a trajectory description, and GPT-4 to provide the final evaluation.
@@
We experiment with evaluating and refining the current state-of-the-art digital agents. In WebArena, this is a GPT-4-based agent described by Zhou et al. (2024). For each task, GPT-4 is provided the user’s instruction and the current DOM representation of the webpage derived from its HTML accessibility tree. GPT-4 is prompted to generate an action grounded in the DOM, e.g., clicking a button with a specific element ID. This agent achieves an end-to-end task success rate of 14.4% using the oracle evaluator.
The strongest agent on the AitW benchmark is CogAgent (Hong et al., 2023), followed by Auto-UI{large, ${}_{\text{\{large, base\}}}$ (Zhang & Zhang, 2023). These agents are implemented as neural vision-language models that map observations, represented as images, and instructions to executable actions. We also experiment with the human demonstrations provided in AitW. [^7] The human demonstrations use the original AitW emulator, which was not released by the authors; thus, these results are not directly comparable with the automated policies, which use the emulator we implement. However, the focus of our experiments is not to directly compare policies, but to compare evaluators across a variety of policies, tasks, and domains.
@@ 4.1 Automatic Evaluation
For each WebArena task and corresponding trajectory sampled from the GPT-4-based policy (Zhou et al., 2024), we acquire task-completion judgments for each of the three evaluation systems described above. Table 1 shows the overall accuracy of the evaluator’s predictions. [^8] Figure 4 in Section A.3 includes the confusion matrices of these predictions. The end-to-end approach with GPT-4V achieves 80.6% accuracy, while Captioner + Mixtral, which uses only open-weight models, matches the oracle’s evaluations for 74.4% of tasks, and replacing Mixtral with GPT-4 achieves the highest accuracy at 82.1%.
GPT-4V Captioner + GPT-4 Captioner + Mixtral
WebArena (%) 80.6 82.1 74.4
Android (%) 90.6 89.8 92.9
@@ 4.2 Autonomous Refinement
Filtered Behavior Cloning on iOS
We demonstrate how our evaluator can guide the refinement of a policy in low-resouce domain using filtered behavior cloning (filtered BC), without additional supervision. We use CogAgent (Hong et al., 2023) as the policy model for the experiment. CogAgent is an 18B-parameter VLM specialized in GUI understanding and navigation. It is primarily instruction-tuned with demonstrations from web navigation and Android device control, and incorporates a very limited, manually collected iOS dataset for training. Given a screenshot and an instruction, CogAgent first generates a high-level plan, followed by the low-level action. For data collection and testing purposes, we design 132 common tasks on iOS, with 80 tasks for training and 52 for testing. Given scaling limitations of emulation, including low speeds and restriction to emulation on macOS, we only experiment with iOS built-in apps and with the Captioner + Mixtral evaluator.
@@
Table 2 contains results for the 52 test tasks. iOS device control is a challenging task, with the baseline agent completing only 8 out of 52 tasks, yielding a 15% success rate. Self-training improves over the baseline by 3 tasks. Filtered BC with our evaluator significantly improved the policy model’s performance from 8 to 14 successes, marking a 75% relative improvement.
Policy # Successful Tasks
CogAgent Baseline 8
@@
Table 2: Comparison of CogAgent and refined policies via self-training and filtered behavior cloning, including the number of successful tasks in our test set (out of 52 total tasks).
@@ 4.3 Error Analysis
Additionally, we find the model provides the correct final evaluation, but incorrect reasoning, for about 10% of correct evaluations.
@@ Appendix A Experiment Details
In this section, we provide details about the implementation of our experiments. Please refer to our code at https://github.com/Berkeley-NLP/Agent-Eval-Refine for the official reference.
@@ A.1 End-to-End Approach
We use gpt-4-1106-vision-preview through the OpenAI API and feed the image without resizing in “high-resolution”. We use a temperature of 0 and keep the other parameters at their default settings. The prompt templates for each environment are provided in Figures 12 and 13.
@@ A.2 Modular Caption-then-Reason Approach
We represent actions as strings, e.g., Type ‘‘Hello’’ . This method leads to information loss when processing actions like clicks for pixel-based policies, as the click coordinates [x, y] become meaningless when the image is represented by its textual description. We leave the task of more adequately transforming pixel-localized actions into textual forms for future work.
@@
The prompt template to query the finetuned Qwen-VL captioner is provided in Figure 16. We fine-tuned the model over 3 epochs with a batch size of 72 and adamw optimizer (Kingma & Ba, 2014), employing a cosine scheduler for learning rate adjustments starting from 1e-5, a weight decay of 0.1, and a warmup ratio of 0.01. As shown in the prompt template, during both finetuning and at inference time, we provide text recognition results from the EasyOCR engine as an additional input to the model to reduce hallucination. We provide randomly-sampled model output examples in Figure 14.
@@ A.4 Evaluation on Android
We use greedy decoding, i.e., temperature of 0, for all policies during evaluation. The confusion matrices comparing our evaluators with human judgments are presented in Figure 4 (right).
@@ A.5 Refinement on WebArena
We implement the Reflexion agent following the original paper (Shinn et al., 2023). The algorithm involves three key components: an Actor, an Evaluator, and a Self-Reflection module. The Actor generates thoughts and actions in text form based on the state observations it receives, where the actions are parsed into executable commands to step the environment for the next observation. The Evaluator assesses the quality of the outputs produced by the Actor. It computes a reward score based on how well the generated trajectory align with the expected outcomes of the task. If the evaluator assesses the task to be failed, the Self-Reflection model will be evoked to generate verbal reflections, which is stored in the agent’s memory and allows the agent to understand its past actions and their outcomes, facilitating improved decision-making in subsequent trials.

### v3

@@ Abstract
We show that domain-general automatic evaluators can significantly improve the performance of agents for web navigation and device control. We experiment with multiple evaluation models that trade off between inference cost, modularity of design, and accuracy. We validate the performance of these models in several popular benchmarks for digital agents, finding between 74.4 and 92.9% agreement with oracle evaluation metrics. Finally, we use these evaluators to improve the performance of existing agents via fine-tuning and inference-time guidance. Without any additional supervision, we improve state-of-the-art performance by 29% on the popular benchmark WebArena, and achieve around 75% relative improvement in device control settings. We release our code and data at https://github.com/Berkeley-NLP/Agent-Eval-Refine.
@@ 1 Introduction
We first evaluate our proposed approach on its ability to match oracle evaluation metrics using WebArena (Zhou et al., 2024) and Android-in-the-Wild (AitW; Rawles et al., 2023), achieving accuracies up to 82.1 and 92.9% respectively. We then show how these evaluation models can be used to refine existing agents through inference-time guidance or during training, without access to any hand-designed evaluation functions or additional demonstration data (Figure 1, right). When integrated as the reward function in Reflexion (Shinn et al., 2023), the evaluator enhances the best-performing GPT-4 WebArena agent’s success rate by up to 29% of relative improvement. Additionally, we evaluate in both Android and iOS device control settings. For iOS, there is no existing benchmark environment or training data. When using our evaluation models to filter sampled trajectories to be used in behavior cloning, we see relative improvements of around 75%.
@@ 2 Related Work
Early modeling of language-conditioned autonomous agents focused on approaches that include semantic parsing (Allen et al., 2007; Xu et al., 2021; Li et al., 2020), reinforcement learning (Branavan et al., 2009; 2010), and imitation learning (Humphreys et al., 2022). The strength of pretrained language and language-and-vision modeling has renewed interest in building language-conditioned digital agents (Zhang & Zhang, 2023; Hong et al., 2024; Zhou et al., 2024; Deng et al., 2023; Wang et al., 2023a; Gur et al., 2024). For example, baseline approaches to WebArena (Zhou et al., 2024) use few-shot prompting with language-only models, representing the environment state and action space with its document object model (DOM). More recent works in building these agents have moved from language-only modeling to vision-language modeling, representing the environment state space as its rendered pixel representation instead of relying on a DOM. Another line of work has applied inference-time techniques to improve model’s performance, for example with inference-time exploration (Zhang et al., 2023), intermediate plan revision (Zhang et al., 2024) and error correction (Wang et al., 2024), and self-critique (Wu et al., 2024) on GPT-4 or GPT-4V. Concurrent to our work, OS-Copilot (Wu et al., 2024) proposes a self-critique component to autonomously refine Mac device control agents, implementing the critic as a LM that reasons about proposed tool implementations and error messages. In contrast to our work, this critic does not evaluate actual agent behavior in the execution environment or used in model training.
@@
Recently, there has been renewed interest in methods for improving policies at training (Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Abdulhai et al., 2023) or inference (Shinn et al., 2023; Yao et al., 2023; Wu et al., 2024) time without human supervision. Approaches like Reflexion (Shinn et al., 2023) assume access to an external evaluation function, leveraging it as the supervision signal to guide policy improvement at inference time. In contrast, we study applications of inference-time autonomous refinement without requiring access to the external evaluation function, and show that using our proposed domain-general evaluation models improves agent success rate by 29%. Meanwhile, methods which bootstrap policies with supervised training and refine them with reinforcement learning have been widely adopted; our proposed evaluators enable this paradigm in open, realistic digital agent scenarios, providing a relative improvement in performance of over 70%. Concurrent to our work, WebVoyager (He et al., 2024) also explores using GPT-4V as an automated proxy for human evaluation of web agents, though is not the primary focus of their work, and neither performs in-depth analysis of the quality of its judgments, nor explores its applicability to improving agents. More broadly, model-based metrics have been highly successful across various domains, such as evaluating image or text quality (Heusel et al., 2017; Zheng et al., 2024), and assessing success of a robot or human activity (Du et al., 2023). Part of our contribution is to provide strong empirical evidence to extend this success to the domain of real-world digital agents.
@@
In parallel, the community has focused on demonstration-based benchmarks that do not include an executable simulation environment, including PIXELHELP (Li et al., 2020), MoTIF (Burns et al., 2022), Mind2Web (Deng et al., 2023), and AitW (Rawles et al., 2023). Notably, Mind2Web and AitW contain over 2K and 715K human trajectories respectively on a wide range of web navigation and device control tasks. Though primarily used for model training (Rawles et al., 2023; Deng et al., 2023; Hong et al., 2024; Zhang & Zhang, 2023), these datasets are also used for evaluating digital agents through reference-based metrics like action matching score. In this setting, an agent is given the prefix of a human demonstration and evaluated on its prediction of the next action to take. However, this metric requires human demonstrations and does not directly reflect agent’s performance in real-world because it does not account for consequences of an agent’s sequential decision process, alternative actions that diverge from the demonstration.
@@ 3.2 Modular Caption-then-Reason Approach
To improve a captioner’s ability to provide detailed, well-formatted descriptions, we collect a dataset of screenshots paired with descriptions, and use it to fine-tune an open-weight VLM. We first acquire screenshots from a variety web and device control domains, then use GPT-4V to provide an initial detailed description for each screenshot. We manually filter out or fix apparent errors in GPT-4V’s output, resulting a total of 1,263 data points. [^3] Table 3 in Section A.2 contains details of data sources and sizes. We use this data to fine-tune the QWen-VL-chat (Bai et al., 2023) model. During both finetuning and at inference time, we provide text recognition results from EasyOCR [^4] https://github.com/JaidedAI/EasyOCR as an additional input to the VLM to reduce hallucination.
@@ 4 Experiments and Results
Our rationale behind experiment design is to cover a broad range of domains and challenges. We use WebArena for both evaluation and inference-time refinement, as its built-in evaluation functions facilitate direct comparison. Android-in-the-Wild (AitW) is chosen for evaluation since it is widely used for training and evaluating Android agents, and is typically evaluated using a reference-based metric instead of task success. Next, we refine a model through filtered behavior cloning on iOS, where data scarcity poses a significant challenge to supervised methods. We further validate the effectiveness of filtered behavior cloning with a similar, larger experiment on Android.
@@
  • QWen-VL-chat: End-to-end approach (Section 3.1) using QWen-VL-chat. This serves as open-weight model’s baseline performance.
  • Captioner + Mixtral: Modular approach (Section 3.2) using a finetuned QWen-VL-chat (Bai et al., 2023) to generate a trajectory description, and Mixtral (Jiang et al., 2024) to provide the final evaluation.
  • Captioner + GPT-4: Modular approach (Section 3.2) using a finetuned QWen-VL-chat to generate a trajectory description, and GPT-4 to provide the final evaluation.
@@
We experiment with evaluating and refining the current state-of-the-art digital agents. In WebArena, this is a GPT-4-based agent described by Zhou et al. (2024). For each task, GPT-4 is provided the user’s instruction and the current DOM representation of the webpage derived from its HTML accessibility tree. GPT-4 is prompted to generate an action grounded in the DOM, e.g., clicking a button with a specific element ID. The task success rate of this agent is 14.4% as reported in Zhou et al. (2024), whereas our reproduction, as shown in Figure 3, indicates a success rate of 15.6%.
The strongest agent on the AitW benchmark is CogAgent (Hong et al., 2024), followed by Auto-UI{large, ${}_{\text{\{large, base\}}}$ (Zhang & Zhang, 2023). These agents are implemented as neural vision-language models that map observations, represented as images, and instructions to executable actions. We also experiment with the human demonstrations provided in AitW. [^7] The human demonstrations use the original AitW emulator, which was not released by the authors; thus, these results are not directly comparable with the automated policies, which use the emulator we implement. However, the focus of our experiments is not to directly compare policies, but to compare evaluators across a variety of policies, tasks, and domains.
@@ 4.1 Automatic Evaluation
For each WebArena task and corresponding trajectory sampled from the GPT-4-based policy (Zhou et al., 2024), we acquire task-completion judgments for each of the three evaluation systems described above. Table 1 shows the overall accuracy of the evaluator’s predictions. [^8] Figure 4 in Section A.3 includes the confusion matrices of these predictions. The end-to-end approach with GPT-4V achieves 80.6% accuracy, while Captioner + Mixtral, which uses only open-weight models, matches the oracle’s evaluations for 74.4% of tasks, and replacing Mixtral with GPT-4 achieves the highest accuracy at 82.1%. Additionally, we see that using QWen-VL-chat end-to-end significantly underperforms the modular, open-weight Captioner + Mixtral approach, confirming the effectiveness of our modular design choice.
GPT-4V QWen-VL-chat Captioner + Mixtral Captioner + GPT-4
WebArena (%) 80.6 68.0 74.4 82.1
Android (%) 90.6 70.2 92.9 89.8
@@ 4.2 Autonomous Refinement
Filtered Behavior Cloning for Device Control
We demonstrate how our evaluator can guide the refinement of a policy using filtered behavior cloning (filtered BC), without additional supervision. We experiment on both iOS and Android environments.
On iOS, we use CogAgent (Hong et al., 2024) as the policy model for the experiment. It is primarily instruction-tuned with demonstrations from web navigation and Android device control, and incorporates a very limited, manually collected iOS dataset for training. For data collection and testing purposes, we design 132 common tasks on iOS, with 80 tasks for training and 52 for testing. Given scaling limitations of emulation, including low speeds and restriction to emulation on macOS, we only experiment with iOS built-in apps and with the Captioner + Mixtral evaluator.
@@
Table 2 contains results for the 52 test tasks. iOS device control is a challenging task, with the baseline agent completing only 8 out of 52 tasks, yielding a 15% success rate. Self-training improves over the baseline by 3 tasks. Filtered BC with our evaluator significantly improves the policy model’s performance from 8 to 14 successes, marking a 75% relative improvement. We also conduct a preliminary analysis of our per-step evaluator’s accuracy. Human annotators agree with our step-wise evaluator on 43 of 50 state-action pairs randomly sampled from the iOS experiments.
However, our iOS experiments require physical Apple computers for emulation, which are slow and not scalable, leading to natural limitations in scope of our evaluation. To address this, we design a similar but larger experiment on Android. Unlike iOS, Android allows us to use scalable emulators on clusters, enabling efficient repetition of experiments. We use AutoUI-base policy for the experiment, selecting 560 tasks for training and 96 for evaluation. The base policy (AutoUI-base) completes 15 tasks, improving to 18.9 $\pm$ 1.0 (averaged over 3 training runs) with self-training. Filtered BC with our step-wise evaluator still significantly improves, succeeding at 26 $\pm$ 0.8 tasks,r esulting in a 73% relative improvement.
Policy # Success (Total: 52)
CogAgent 8
@@
(a) iOS Experiment
Policy # Success (Total: 96)
Auto-UI-base 15
+ Self-training 18.9 $\pm$ 1.0
+ Filtered BC (Ours) 26.0 $\pm$ 0.8
(b) Android Experiment
Table 2: Comparison of base policy and refined policies via self-training and filtered behavior cloning, including the number of successful tasks in our test set.
@@ 4.3 Error Analysis
Additionally, we find the model provides the correct final evaluation, but incorrect reasoning, in about 10% of correct evaluations.
@@ Appendix A Experiment Details
In this section, we provide details about our experiments. Please refer to our code at https://github.com/Berkeley-NLP/Agent-Eval-Refine for the official reference.
@@ A.1 End-to-End Approach
We use gpt-4-1106-vision-preview through the OpenAI API and feed the image without resizing in “high-resolution”. We use a temperature of 0 and keep other parameters at default. The prompt templates for each environment are provided in Figures 12 and 13.
@@ A.2 Modular Caption-then-Reason Approach
We represent actions as strings, e.g., Type ‘‘Hello’’ . This method leads to information loss when processing actions like clicks for pixel-based policies, as the coordinates [x, y] become meaningless when the image is represented by its textual description. We leave the task of more adequately transforming pixel-localized actions into textual forms for future work.
@@
The prompt template to query the finetuned Qwen-VL-chat captioner is provided in Figure 16. We fine-tuned the model over 3 epochs with a batch size of 72 and adamw optimizer (Kingma & Ba, 2014), employing a cosine scheduler for learning rate adjustments starting from 1e-5, a weight decay of 0.1, and a warmup ratio of 0.01. As shown in the prompt template, during both finetuning and at inference time, we provide text recognition results from the EasyOCR engine as an additional input to the model to reduce hallucination. We provide randomly-sampled model output examples in Figure 14.
@@ A.4 Evaluation on Android
We use temperature of 0, for all policies during evaluation. The confusion matrices comparing our evaluators with human judgments are presented in Figure 4 (right).
@@ A.5 Refinement on WebArena
We implement the Reflexion agent following the original paper (Shinn et al., 2023). The algorithm involves three key components: an Actor, an Evaluator, and a Self-Reflection module. The Actor generates thoughts and actions in text based on the state observations it receives, where the actions are parsed into executable commands to transform the environment. The Evaluator assesses the quality of the outcome produced by the Actor. It computes a reward score based on how well the generated trajectory align with the goal. If the evaluator assesses the task to be failed, the Self-Reflection model will be evoked to generate verbal reflections, which is stored in the actor’s memory, facilitating improved decision-making in subsequent trials.